The Candy Company sells candy, nuts and raisins. They package their nuts and raisins together in 2 different ratio packs. One of their packs is 50% raisins and 50% nuts. The other mix is 25% raisins and 75% nuts. The price on the 50-50 mix is $7. The price on the 25-75 mix is $9.50. The current situation is that there are 114 kg of nuts and 82 kg of raisins. How should they package their nuts and raisins to maximize the revenue? Set up an Excel spreadsheet to maximize the revenue using solver and the constraints listed.
Answer #1

We have x => 0 and y => 0

Revenue = 7x + 9.5y

Mix A has equal combination of nuts and raisin

1/2x + 3/4y <= 114

For mix b

1/2x + 1/4y <= 82

2x + 3y <= 456 and

2x + y <= 328

By excel solver,

So,

Nuts should be 132 kg and raisins should be 64 kg to provide a maximum revenue of $1532.

please leave a like. Thank you.